# Efficient 3D Molecular Generation with Flow Matching and Scale Optimal Transport

### v2

@@ 3.3 Overall Architecture
As shown in Fig. 2 a full Semla model consists of a stack of $L$ Semla layers along with embedding layers and MLPs for encoding the atom and bond types, and MLP prediction heads for producing unnormalised distributions for atoms, bonds and formal charges. Unlike EQGAT, our model does not carry edge features throughout the network. Instead, the first layer embeds bond information into the node features by passing the encoded bond features into the pairwise message module. Analogously, the final layer produces pairwise edge features which are then further updated through a bond refinement layer at the end of the network. This layer acts in a similar way to the pairwise message block described above but only updates the edge features. Absorbing bond information into the node features like this leads to a further increase in the efficiency of our model and, in our experiments, had little impact on generative performance.
@@ 4 Flow Matching for Molecular Generation
To assess the ability of Semla to model distributions with $E(3)\times S_{N}$ symmetry we apply conditional flow matching [28, 2, 1, 30] with optimal transport to create a generative model for molecules, which we refer to as SemlaFlow. In this section we outline the training and sampling procedure for SemlaFlow with molecular structures. In the following we represent each molecule by a tuple $z=(\mathbf{x},\mathbf{a},\mathbf{b},\mathbf{c})$ of coordinates, atom types, bond types and formal charges, respectively.
@@
As shown in existing conditional flow matching frameworks [28, 2, 10], training proceeds by firstly sampling: noise $z_{0}\sim p_{\mathrm{noise}}(z_{0})$; data $z_{1}\sim p_{\mathrm{data}}(z_{1})$; and a time $t\in\left[0,1\right]$, and using these to sample from the time-dependent conditional flow $z_{t}\sim p_{t|1}(z|z_{0},z_{1})$. In practice, we apply scale optimal transport (OT) and sample noise from the scaled distribution $p_{\mathrm{noise}}^{\pi}(z_{0}|n)$, where $n$ is the number of atoms. We discuss this further, along with the OT transformation applied to the noise during training below. Previous work on molecular structure generation has found it beneficial to train models to predict data directly rather than noise [26] or a vector field [48]. SemlaFlow is therefore trained to learn a distribution $p_{1|t}^{\theta}(z_{1}|z_{t})$ which predicts clean data from interpolated data using a Semla model with parameters $\theta$. Different conditional flows are used for continuous and discrete data. For coordinates, we apply a linear interpolation with the addition of a small amount of Gaussian noise. Atom and bond types are sampled according to the recently proposed discrete flow models (DFM) framework [10]. The joint molecular interpolation is therefore given as follows:
$\mathbf{x}_{t}\sim\mathcal{N}(t\mathbf{x}_{1}-(1-t)\mathbf{x_{0}},\sigma^{2})\hskip 25.0ptt\sim\text{Beta}(\alpha,\beta)$ (10)
$a_{t}\sim\text{Cat}(t\delta(a_{1})+(1-t)\frac{1}{|\mathcal{A}|})\hskip 25.0ptb_{t}\sim\text{Cat}(t\delta(a_{1})+(1-t)\frac{1}{|\mathcal{B}|})$ (11)
Where $\mathcal{A}$ and $\mathcal{B}$ are the sets of atom and bond types, respectively, and $\delta(\cdot)$ is the one-hot encoding operation. We use $(\alpha,\beta)$ = (2.0, 1.0) and $\sigma=0.2$ for all models presented in this paper. Notably, although our model predicts the formal charge for each atom, the charges do not participate in the generative flow-matching process and so there is no need to interpolate them.
After sampling a predicted molecule ${\tilde{z}}_{1}\sim p_{1|t}^{\theta}(z_{1}|z_{t})$ where ${\tilde{z}}_{1}=(\mathbf{\tilde{x}}_{1},\mathbf{\tilde{a}}_{1},\mathbf{\tilde{b}}_{1},\mathbf{\tilde{c}}_{1})$, the model is trained with a mean-squared error loss function ($\mathcal{L}_{\mathrm{MSE}}$) for coordinates, and cross-entropy losses ($\mathcal{L}_{\mathrm{CE}}$) for atom types, bond types and charges. The final loss for the model is then given by the weighted sum:
$\mathcal{L}_{\mathrm{SemlaFlow}}=\lambda_{x}\mathcal{L}_{\mathrm{MSE}}(\mathbf{\tilde{x}}_{1},\mathbf{x}_{1})+\lambda_{a}\mathcal{L}_{\mathrm{CE}}(\mathbf{\tilde{a}}_{1},\mathbf{a}_{1})+\lambda_{b}\mathcal{L}_{\mathrm{CE}}(\mathbf{\tilde{b}}_{1},\mathbf{b}_{1})+\lambda_{c}\mathcal{L}_{\mathrm{CE}}(\mathbf{\tilde{c}}_{1},\mathbf{c}_{1})$ (12)
We also make use of self-conditioning, which was originally proposed for diffusion models as way of reusing the model’s previous prediction when sampling [11]. To create a self-conditioned SemlaFlow model, we adopt the same training procedure as HarmonicFlow [48]. We provide further details on this, along with the hyperparameters used for SemlaFlow, in Appendix C.
Sampling Molecules
Once we have trained a SemlaFlow model, molecules can be generated by, firstly, sampling noise $z_{0}\sim p_{\mathrm{noise}}^{\pi}(z_{0}|n)$, and then integrating the ODE corresponding to the conditional flow $p_{t|1}$ from $t=0$ to $t=1$. For coordinates, the vector field corresponding to our choice of conditional flow is given by $\mathbf{\tilde{x}}_{1}-\mathbf{x}_{0}$ where $\mathbf{\tilde{x}}_{1}$ is sampled from $p_{1|t}^{\theta}$ as shown above. We then apply an Euler solver to integrate the ODE with step sizes $\Delta t$ as follows: $\mathbf{x}_{t+\Delta t}=\mathbf{x}_{t}+\Delta t(\mathbf{\tilde{x}}_{1}-\mathbf{x}_{0})$. We refer readers to DFM [10] for the sampling procedure for atom and bond types. In practice, we found that taking logarithmically spaced steps resulted in better performance than using constant step sizes.
Scale Optimal Transport
For molecular distributions the average length between atom coordinates and the centre of the molecules increases with the number of atoms $N$. In polymer theory the standard deviation of atom coordinates within self-avoiding polymer chains has been shown to scale with $\sqrt{N}$, known as the Flory radius [39]. Typically, however, for molecular generation models, coordinate noise is sampled from a unit Gaussian which doesn’t reflect differences in molecular sizes. Instead we sample from a Gaussian distribution with a variance that depends on $N$, the number of atoms in $\mathbf{x}_{1}$. We produce a sample $\mathbf{x}_{0}^{\pi}$ from the noise distribution $p_{\mathrm{noise}}^{\pi}$ during training and inference as follows:
$\mathbf{x}_{0}^{\pi}=\begin{cases}f_{\pi}(\mathbf{x}_{0},\mathbf{x}_{1})&\hskip 10.0pt\text{training}\\ \mathbf{x}_{0}&\hskip 10.0pt\text{inference}\end{cases}\hskip 40.0pt\mathbf{x}_{0}\sim\mathcal{N}(0,\sigma_{N}^{2})$ (13)
Where $f_{\pi}(\mathbf{x}_{0},\mathbf{x}_{1})$ is the equivariant optimal transport [23, 47] transformation which corresponds to applying a permutation and rotation which minimises the transport cost (in this case the mean-squared error) between $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$. Since existing 3D generative models already require $N$ to be known at inference time, no additional restrictions are placed on the sampling when using scale OT.
Since it applies to long polymer chains, we consider the Flory radius as an upper-bound for small molecule drugs, and empirically find $\sigma_{N}=k\log(N)$ with $k=0.2$ to be a good fit for a dataset of drug-like molecules. In Appendix A we provide further details on this approach and show that it leads to a significant reduction in the transport cost compared to equivariant optimal transport, especially for smaller molecules.
5 Experiments
In this section we provide results on benchmark 3D molecular generation tasks and compare the performance of our model to existing state-of-the-art approaches. In Appendix B, we provide results on ablation experiments which show how different components of the training, such as scale OT and self conditioning, affect performance on molecular generation tasks. We also provide samples from a SemlaFlow model trained on GEOM Drugs in Appendix E.
Datasets
Two benchmark datasets, QM9 [37] and GEOM Drugs [4], are used to assess SemlaFlow’s abilities as an unconditional molecular generator. QM9 is a complete enumeration of molecules with up to 9 heavy atoms which satisfy a set of pre-defined constraints. On average, molecules in QM9 contain approximately 18 atoms. GEOM Drugs, on the other hand, is a dataset of larger, drug-like molecules with an average size of around 44 atoms. For both datasets we use the same data splits as MiDi and EQGAT-diff. To improve training times, however, we discard molecules with more than 72 atoms from the GEOM Drugs training set. This corresponds to about 1% of the training data. Validation and test sets are left unchanged. All metrics for SemlaFlow presented below are calculated by, firstly, sampling molecule sizes from the test set, and then generating molecules with the sampled number of atoms using the model, as outlined above.
@@ 5 Experiments
Table 1: Molecular generation results on QM9. Values for models we evaluated are given as an average over 5 runs, sampling 5000 molecules on each run, with standard deviations given in subscripts. Models are grouped into those which infer bonds from coordinates (top) and those which generate bonds directly (bottom). Since some models only publish the proportion of molecules which are both unique and valid, results marked are estimates for uniqueness.
Model Atom Stab $\uparrow$ Mol Stab $\uparrow$ Valid $\uparrow$ Unique $\uparrow$ NFE $\downarrow$
@@
SemlaFlow (Ours) 99.9±${}_{\pm\text{0.0}}$ 99.6±${}_{\pm\text{0.06}}$ 99.4±${}_{\pm\text{0.12}}$ 97.5±${}_{\pm\text{0.20}}$ 100
Baselines
@@
We compare SemlaFlow to a number of recently proposed models for 3D molecular generation, including both models which infer bonds based on coordinates and models which generate bonds directly. EDM [19], GCDM [33], MUDiff [20] and GFMDiff [58] are all diffusion-based models which infer bonds from atom positions. We also compare to EquiFM [47] which uses flow-matching along with equivariant optimal transport to generate atom types and coordinates; bonds are then inferred based on these. Finally, we also compare to MiDi [56] and EQGAT-diff [26] which learn molecular bonds and which we regard as the current state-of-the-art. We use standard benchmark evaluation metrics: atom stability; molecule stability; validity; uniqueness; and novelty, which have been thoroughly described in previous works. We also provide a full description of these metrics in Appendix D. Additionally, for each baseline model we provide the number of function evaluations (NFE) required to sample one batch of molecules. These were taken from the respective publications, if available, or using the default values in the official codebases.
@@
Table 1 compares the performance of SemlaFlow with existing approaches. Following [55, 19] we do not provide novelty scores on QM9. SemlaFlow is trained for 300 epochs on QM9 on a single Nvidia A100 GPU, and we provide further training details in Appendix C. From the table we can see that SemlaFlow matches or exceeds all models on all metrics other than uniqueness, despite using 5 times fewer sampling steps than MiDi and EQGAT-diff. Our model also significantly outperforms EquiFM, the only other flow-matching-based model in the table. We also note that since EquiFM uses an adaptive ODE solver the number of function evaluations in the table is an average and could be significantly higher for larger, more complex molecules.
@@
Since QM9 contains only very small molecules GEOM Drugs serves a more useful benchmark for distinguishing model performance. Table 2 compares our model to MiDi and EQGAT-diff on this dataset. We do not include models which infer bonds since these models typically perform very poorly on larger molecules and often do not provide results for all evaluation metrics. Once again we see that SemlaFlow outperforms existing state-of-the-art models on all key evaluation metrics, while generating only marginally fewer novel molecules. In addition to requiring significantly fewer evaluation steps our model also requires much less compute for training. SemlaFlow trains for 200 epochs on a single Nvidia A100 GPU, compared to 800 epochs with 4 GPUs for EQGAT-diff.
Table 2: Molecular generation results on GEOM Drugs. Values for models we evaluated are given as an average over 5 runs, sampling 5000 molecules on each run, with standard deviations in subscripts.
Model Atom Stab $\uparrow$ Mol Stab $\uparrow$ Valid $\uparrow$ Unique $\uparrow$ Novel $\uparrow$ NFE $\downarrow$
MiDi 99.8 91.6 77.8 100.0 100.0 500
EQGAT-diff 99.8±${}_{\pm\text{0.01}}$ 93.4±${}_{\pm\text{0.21}}$ 94.6±${}_{\pm\text{0.24}}$ 100.0±${}_{\pm\text{0.0}}$ 99.9±${}_{\pm\text{0.07}}$ 500
SemlaFlow (Ours) 99.8±${}_{\pm\text{0.01}}$ 97.7±${}_{\pm\text{0.07}}$ 95.2±${}_{\pm\text{0.22}}$ 100.0±${}_{\pm\text{0.0}}$ 99.6±${}_{\pm\text{0.09}}$ 100
Sampling Efficiency
To further improve the efficiency of our model we tried varying the number of ODE integration steps and we provide a one-to-one comparison of these results with EQGAT-diff in Table 3. We also measure the time required by each model to sample 5000 molecules. For EQGAT-diff we measure the sample time using the given evaluation code with a batch size of 75, which was the largest multiple of 25 we could consistently fit on an Nvidia A100 40GB GPU. In the table we can see a stark difference in runtime between the two models. SemlaFlow provides comparable performance to EQGAT-diff with as few as 20 sampling steps, corresponding to a two order of magnitude speed-up in sampling time. Much of this improvement is due to the very small number of sampling steps, which we hypothesise is due to the use of optimal transport, since this aims to minimise the cost of transporting probability mass between distributions. However, we further see a two-fold improvement in the time per forward pass, which demonstrates the importance of designing efficient equivariant architectures for molecular generation.
Conformer Evaluation
The validity and stability metrics presented here, however, only measure the 1D structure of the molecule (i.e. the SMILES string); they provide no information on the quality of the conformations. We can also see that recent models have saturated these metrics on both QM9 and GEOM Drugs, warranting other evaluation methods. In Appendix D we outline additional problems with current metrics. To further compare model performance and to allow evaluation of 3D generation, we introduce energy and strain energy as new benchmark metrics for this task. The energy measures the quality of a conformer, considering typical bonded and non-bonded interactions. The energy $U(\mathbf{x})$ of a conformation is inversely related to its probability according to the Boltzmann distribution $p(\mathbf{x})=Z^{-1}\exp\left(-U(\mathbf{x})/kT\right)$ where $T$ is the temperature and $k$ is the Boltzmann constant. The strain is given by the difference $U(\mathbf{x})-U(\mathbf{\tilde{x}})$ where $\mathbf{\tilde{x}}$ is the relaxed (i.e. minimised) conformation for $\mathbf{x}$. We argue that these metrics provide a useful overview of the quality of the generated conformations and directly include measurements such as bond lengths and bond angles which have been proposed previously [56]. We use RDKit [14] with an MMFF94 [16] forcefield to calculate the energies and perform the minimisation.
@@
Table 3: Comparison between EQGAT-diff and SemlaFlow with different numbers of sampling steps. Molecule stability and validity are given as percentages, energy and strain energy are measured in $\mathrm{kcal}\cdot\mathrm{mol}^{-1}$, NFE refers to the number of function evaluations required to sample one batch of molecules and sample time is measured by the number of seconds required to generate 5000 molecules, averaged over 5 runs.
@@
Model Mol Stab $\uparrow$ Valid $\uparrow$ Energy $\downarrow$ Strain $\downarrow$ NFE $\downarrow$ Sample Time $\downarrow$
EQGAT-diff 93.4±${}_{\pm\text{0.21}}$ 94.6±${}_{\pm\text{0.24}}$ 148.8±${}_{\pm\text{0.88}}$ 140.2±${}_{\pm\text{0.72}}$ 500 11521
SemlaFlow20 94.8±${}_{\pm\text{0.28}}$ 93.7±${}_{\pm\text{0.36}}$ 168.7±${}_{\pm\text{3.1}}$ ${}_{\pm\text{3.1}}$ 20 167
SemlaFlow50 97.2±${}_{\pm\text{0.25}}$ 95.2±${}_{\pm\text{0.40}}$ 141.1±${}_{\pm\text{2.1}}$ 102.0±${}_{\pm\text{2.1}}$ 50 412
SemlaFlow100 97.7±${}_{\pm\text{0.07}}$ 95.2±${}_{\pm\text{0.22}}$ ${}_{\pm\text{2.1}}$ 88.9±${}_{\pm\text{1.3}}$ 100 821
Data 100.0 100.0 50.3 15.9
@@
Table 3 shows that SemlaFlow produces molecules with better energies and strain energies than EQGAT-diff, despite the significantly faster sampling time. Notably, however, molecules generated by EQGAT-diff have lower minimised energies than SemlaFlow, suggesting that their model is better at finding molecular conformations which have lower energy minima, while our model is better at producing lower strain energies. In ablation experiments in Appendix B we also show that the use of scale optimal transport is a crucial component for generating low energy conformations with our model.

### v3

@@ 3.3 Overall Architecture
A full Semla model consists of a stack of Semla layers along with embedding layers and MLPs for encoding the atom and bond types, and MLP prediction heads for producing unnormalised distributions for atoms, bonds and formal charges. Unlike EQGAT, our model does not carry edge features throughout the network. Instead, the first layer embeds bond information into the node features by passing the encoded bond features into the pairwise message module. Analogously, the final layer produces pairwise edge features which are then further updated through a bond refinement layer at the end of the network. This layer acts in a similar way to the pairwise message block described above but only updates the edge features. Absorbing bond information into the node features like this leads to a further increase in the efficiency of our model and, in our experiments, had little impact on generative performance.
Unless stated otherwise, the Semla models we present in the remainder of this paper are constructed from 12 layers with $d_{inv}=384$, $d_{equi}=64$, $d_{l}=64$, $n_{heads}=32$. This corresponds to approximately 22M learnable parameters. We provide a full model overview and further hyperparameter details in Appendix C.
@@ 4 FLOW MATCHING FOR MOLECULAR GENERATION
To assess the ability of Semla to model distributions with $E(3)\times S_{N}$ symmetry we apply conditional flow matching (Lipman et al., 2023; Albergo and Vanden-Eijnden, 2023; Albergo et al., 2023; Liu et al., 2023) with optimal transport to create a generative model for molecules, which we refer to as SemlaFlow. Unlike many previous approaches to 3D generation, we learn to generate a join distribution over atomic coordinates, atom types, bond orders and formal charges, rather than inferring parts of this distribution after generation.
Each molecule is represented by a tuple $z=(\mathbf{x},\mathbf{a},\mathbf{b},\mathbf{c})$ of coordinates, atom types, bond orders and formal charges, respectively. Since we wish to train a model to sample from the joint distribution $p(z)$ which contains both discrete and continuous data, we parameterise a single Semla model to generate multiple vector fields. For discrete data (atom and bond types) we apply the discrete flow models (DFM) framework (Campbell et al., 2024), and for the continuous atom coordinates we apply the flow matching algorithm proposed by Tong et al. (2024), where a small amount of gaussian noise is added the interpolated coordinates $x_{t}$. Our model also predicts the formal charge for each atom but these do not participate in the generative flow matching process. In the remainder of this section we outline the full training and sampling procedure for SemlaFlow with molecular structures.
@@
As shown in existing conditional flow matching frameworks (Lipman et al., 2023; Albergo and Vanden-Eijnden, 2023; Campbell et al., 2024), training proceeds by firstly sampling: noise $z_{0}\sim p_{\mathrm{noise}}(z_{0})$; data $z_{1}\sim p_{\mathrm{data}}(z_{1})$; and a time $t\in\left[0,1\right]$, and using these to sample from the time-dependent conditional flow $z_{t}\sim p_{t|1}(z|z_{0},z_{1})$. The joint molecular interpolation is therefore given as follows:
$\mathbf{x}_{t}\sim\mathcal{N}(t\mathbf{x}_{1}+(1-t)\mathbf{x_{0}},\sigma^{2})\hskip 15.0ptt\sim\text{Beta}(\alpha,\beta)$ (14)
$a_{t}\sim\text{Cat}(t\delta\{a_{1},a_{t}\}+(1-t)\frac{1}{|\mathcal{A}|})$ (15)
$b_{t}\sim\text{Cat}(t\delta\{b_{1},b_{t}\}+(1-t)\frac{1}{|\mathcal{B}|})$ (16)
Where $\mathcal{A}$ and $\mathcal{B}$ are the sets of atom and bond types, respectively, and $\delta\{i,j\}$ is the Kronecker delta which is $1$ when $i=j$ and 00 otherwise. We use $(\alpha,\beta)=(2.0,1.0)$ and $\sigma=0.2$ for all SemlaFlow models presented in this paper. In practice, we apply the equivariant optimal transport (OT) (Klein et al., 2024; Song et al., 2023) transformation to the sampled coordinates $\hat{\mathbf{x}}_{0}=f_{\pi}(\mathbf{x}_{0},\mathbf{x}_{1})$ before sampling the interpolated value $\mathbf{x}_{t}$. This corresponds to applying a permutation and rotation which minimises the transport cost (in this case the mean-squared error) between $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$.
@@
Table 1: Molecular generation results on QM9. Models are grouped into those which infer bonds from coordinates (top) and those which generate bonds directly (bottom). Since some models only publish the proportion of molecules which are both unique and valid, results marked are estimates for uniqueness.
Model Atom Stab $\uparrow$ Mol Stab $\uparrow$ Valid $\uparrow$ Unique $\uparrow$ NFE
@@
SemlaFlow (Ours) 99.9±${}_{\pm\text{0.0}}$ 99.7±${}_{\pm\text{0.03}}$ 99.4±${}_{\pm\text{0.03}}$ 95.4±${}_{\pm\text{0.12}}$ 100
Previous work on molecular structure generation has found it beneficial to train models to predict data directly rather than noise (Le et al., 2024) or a vector field (Stark et al., 2024a). SemlaFlow is therefore trained to learn a distribution $p_{1|t}^{\theta}(z_{1}|z_{t})$ which predicts clean data from interpolated data using a Semla model with parameters $\theta$. After sampling a predicted molecule ${\tilde{z}}_{1}\sim p_{1|t}^{\theta}(z_{1}|z_{t})$ where ${\tilde{z}}_{1}=(\mathbf{\tilde{x}}_{1},\mathbf{\tilde{a}}_{1},\mathbf{\tilde{b}}_{1},\mathbf{\tilde{c}}_{1})$, the model is trained with a mean-squared error loss function ($\mathcal{L}_{\mathrm{MSE}}$) for coordinates, and cross-entropy losses ($\mathcal{L}_{\mathrm{CE}}$) for atom types, bond types and charges. The final loss for the model is then given by the weighted sum:
$\mathcal{L}_{\mathrm{SemlaFlow}}=\lambda_{x}\mathcal{L}_{\mathrm{MSE}}(\mathbf{\tilde{x}}_{1},\mathbf{x}_{1})+\lambda_{a}\mathcal{L}_{\mathrm{CE}}(\mathbf{\tilde{a}}_{1},\mathbf{a}_{1})$ (17)
$+\lambda_{b}\mathcal{L}_{\mathrm{CE}}(\mathbf{\tilde{b}}_{1},\mathbf{b}_{1})+\lambda_{c}\mathcal{L}_{\mathrm{CE}}(\mathbf{\tilde{c}}_{1},\mathbf{c}_{1})$ (18)
We also make use of self-conditioning, which was originally proposed for diffusion models as way of reusing the model’s previous prediction when sampling (Chen et al., 2023). To create a self-conditioned SemlaFlow model, we adopt the same training procedure as HarmonicFlow (Stark et al., 2024a). We provide further details on this, along with the hyperparameters used for SemlaFlow, in Appendix C.
Sampling Molecules
Once we have trained a SemlaFlow model, molecules can be generated by, firstly, sampling noise $z_{0}\sim p_{\mathrm{noise}}(z_{0}|n)$, and then integrating the ODE corresponding to the conditional flow $p_{t|1}$ from $t=0$ to $t=1$. For coordinates, the vector field corresponding to our choice of conditional flow is given by $\mathbf{\tilde{x}}_{1}-\mathbf{x}_{0}=\frac{1}{1-t}(\mathbf{\tilde{x}}_{1}-\mathbf{x}_{t})$ where $\mathbf{\tilde{x}}_{1}\sim p_{1|t}^{\theta}$ as shown above. We then apply an Euler solver to integrate the ODE with step sizes $\Delta t$ as follows: $\mathbf{x}_{t+\Delta t}=\mathbf{x}_{t}+\frac{\Delta t}{1-t}(\mathbf{\tilde{x}}_{1}-\mathbf{x}_{t})$. We refer readers to DFM (Campbell et al., 2024) for the sampling procedure for atom and bond types. In practice, we found that taking logarithmically spaced steps, where the model spends more time in parts of the vector field closer to $t=1$, resulted in better performance than using constant step sizes.
5 EXPERIMENTS
In this section we provide results on benchmark 3D molecular generation tasks and compare the performance of our model to existing state-of-the-art approaches. In Appendix A we also provide results on multiple ablation experiments which demonstrate the impact of various contributions of this paper. One experiment demonstrates how controlling the latent message size $d_{l}$ can lead to a significant model speedup without impacting performance. Another provides a side-by-side comparison of the Semla architecture with the EQGAT and EGNN architectures trained using identical flow matching training setups. We show that Semla outperforms the EQGAT and EGNN architectures at various model sizes, while also providing between 3 and 5 times faster inference over larger models. Finally, we provide samples from a SemlaFlow model trained on GEOM Drugs in Appendix E.
Evaluation Setup
Two benchmark datasets, QM9 (Ramakrishnan et al., 2014) and GEOM Drugs (Axelrod and Gomez-Bombarelli, 2022), are used to assess SemlaFlow’s abilities as an unconditional molecular generator. Since QM9 contains only very small molecules GEOM Drugs serves a more useful benchmark for distinguishing model performance. For both datasets we use the same data splits as MiDi and EQGAT-diff. To improve training times, however, we discard molecules with more than 72 atoms from the GEOM Drugs training set. This corresponds to about 1% of the training data; validation and test sets are left unchanged. All metrics for SemlaFlow presented below are calculated by sampling from the distribution of molecule sizes in the test set, and then generating molecules with the sampled number of atoms by integrating the trained ODE.
Table 2: Molecular generation results on GEOM Drugs. Results for models which perform bond inference are provided in the appendix since they often do not provide results for all standard benchmark metrics.
Model Atom Stab $\uparrow$ Mol Stab $\uparrow$ Valid $\uparrow$ Unique $\uparrow$ Novel $\uparrow$ NFE
FlowMol 99.0 67.5 51.2 100
MiDi 99.8 91.6 77.8 100.0 100.0 500
EQGAT-diff 99.8±${}_{\pm\text{0.0}}$ 93.4±${}_{\pm\text{0.21}}$ 94.6±${}_{\pm\text{0.24}}$ 100.0±${}_{\pm\text{0.0}}$ 99.9±${}_{\pm\text{0.07}}$ 500
SemlaFlow (Ours) 99.8±${}_{\pm\text{0.0}}$ 97.3±${}_{\pm\text{0.08}}$ 93.9±${}_{\pm\text{0.19}}$ 100.0±${}_{\pm\text{0.0}}$ 99.6±${}_{\pm\text{0.03}}$ 100
@@ 5 EXPERIMENTS
We compare SemlaFlow to a number of recently proposed models for 3D molecular generation. EDM (Hoogeboom et al., 2022), GCDM (Morehead and Cheng, 2024), GFMDiff (Xu et al., 2024), MUDiff (Hua et al., 2024), GeoLDM (Xu et al., 2023) and GeoBFN (Song et al., 2024) are all diffusion-based models which infer bonds from atom positions. We also compare to EquiFM (Song et al., 2023) which uses flow-matching along with equivariant optimal transport to generate atom types and coordinates; bonds are then inferred based on these. FlowMol (Dunn and Koes, 2024) is a recently proposed flow matching model which learns a joint distribution over atoms and bonds and is probably the most similar existing work to ours. Finally, we also compare to MiDi (Vignac et al., 2023) and EQGAT-diff (Le et al., 2024) which we regard as the existing state-of-the-art. We use standard benchmark evaluation metrics: atom stability; molecule stability; validity; uniqueness; and novelty, which have been thoroughly described in previous works. We also provide a full description of these metrics in Appendix D. Results for models we evaluated are given as an average over 3 runs, sampling 10,000 molecules on each run, with standard deviations provided in subscripts. We also provide the number of function evaluations (NFE) required to sample one batch of molecules.
@@
Table 1 compares the performance of SemlaFlow with existing approaches on the QM9 dataset. Following Vignac and Frossard (2022); Hoogeboom et al. (2022) we do not provide novelty scores on QM9. SemlaFlow is trained for 300 epochs on QM9 on a single Nvidia A100 GPU. From the table we can see that SemlaFlow matches or exceeds all models on all metrics other than uniqueness, despite using 5 times fewer sampling steps than MiDi and EQGAT-diff. Our model also outperforms EquiFM and FlowMol, the only other flow matching models in the table.
@@
Table 2 compares SemlaFlow’s performance on GEOM Drugs to existing models which are also trained to generate molecular bonds. SemlaFlow matches or exceeds the performance of existing models on atom and molecule stability and produces only slightly fewer valid molecules than EQGAT-diff. The performance difference between FlowMol and SemlaFlow is also much more noticeable than on QM9; only two-thirds of molecules produced by FlowMol satisfy the molecular stability metric, compared to more than 97% for SemlaFlow. In addition to requiring significantly fewer evaluation steps than MiDi and EQGAT-diff, our model also requires much less compute for training. SemlaFlow trains for 200 epochs on a single Nvidia A100 GPU, compared to 800 epochs with 4 GPUs for EQGAT-diff.
Due to space constraints and since these models often do not provide results for all standard metrics, we provide results for models which infer bonds in Appendix B. Notably, while some of these models have higher validity than SemlaFlow, most have molecular stabilities close to 0 – they struggle to generate molecules with the correct number of bonds for each atom, especially on larger, drug-like molecules, like those in GEOM Drugs. Since the validity metric only measures whether a molecule can be loaded into RDKit (Gred Landrum et al., 2023), we consider molecule stability a more comprehensive metric for this task. We discuss this issue further, along with other problems with existing evaluation metrics, in Appendix D.
Further Evaluation
@@
Table 3: Comparison between EQGAT-diff and SemlaFlow with different numbers of sampling steps. Energy and strain energy are given as an average per atom and are measured in $\mathrm{kcal}\cdot\mathrm{mol}^{-1}$. Sample time is measured by the average number of seconds to generate 1000 molecules. All metrics are averaged over 3 runs.
@@
Model Mol Stab $\uparrow$ Valid $\uparrow$ Energy $\downarrow$ Strain $\downarrow$ Sample Time $\downarrow$ NFE
EQGAT-diff 93.4±${}_{\pm\text{0.21}}$ 94.6±${}_{\pm\text{0.24}}$ 3.38±${}_{\pm\text{0.020}}$ 3.23±${}_{\pm\text{0.020}}$ 2293.0 500
SemlaFlow20 95.3±${}_{\pm\text{0.14}}$ 93.0±${}_{\pm\text{0.10}}$ 2.58±${}_{\pm\text{0.018}}$ 1.76±${}_{\pm\text{0.007}}$ 20.3 20
SemlaFlow50 97.0±${}_{\pm\text{0.21}}$ 93.9±${}_{\pm\text{0.12}}$ 2.33±${}_{\pm\text{0.044}}$ 1.46±${}_{\pm\text{0.043}}$ 49.8 50
SemlaFlow100 97.3±${}_{\pm\text{0.08}}$ 93.9±${}_{\pm\text{0.19}}$ 2.24±${}_{\pm\text{0.028}}$ 1.37±${}_{\pm\text{0.022}}$ 99.3 100
Data 100.0 100.0 1.12 0.36
The validity and stability metrics presented here, however, only measure the topological structure of the molecule (i.e. the molecular graph); they provide no information on the quality of the generated conformations. We can also see that some existing models have already saturated these metrics on both QM9 and GEOM Drugs, warranting the introduction of additional evaluation methods. To further compare model performance and to allow evaluation of 3D generation, we introduce energy per atom and strain energy per atom as new benchmark metrics for this task. The energy measures the quality of a conformer, considering typical bonded and non-bonded interactions. The energy $U(\mathbf{x})$ of a conformation is inversely related to its probability according to the Boltzmann distribution $p(\mathbf{x})=Z^{-1}\exp\left(-U(\mathbf{x})/kT\right)$ where $T$ is the temperature and $k$ is the Boltzmann constant. The strain is given by the difference $U(\mathbf{x})-U(\mathbf{\tilde{x}})$ where $\mathbf{\tilde{x}}$ is the relaxed (i.e. minimised) conformation for $\mathbf{x}$. Since molecular energy is generally calculated as a sum of atomic energies we normalise both metrics by the number of atoms in each generated molecule. We argue that these metrics provide a very useful overview of the quality of the generated conformations and directly include measurements such as bond lengths and bond angles which have been proposed previously (Vignac et al., 2023; Buttenschoen et al., 2024). We use RDKit (Gred Landrum et al., 2023) with an MMFF94 (Halgren, 1996) forcefield to calculate the energies and perform the minimisation.
@@
In Table 3 we provide a performance and sampling time comparison between EQGAT-diff and SemlaFlow with varying numbers of ODE integration steps. SemlaFlow produces higher molecule stabilities and better energies and strain energies than EQGAT-diff, even with as few as 20 integration steps, although the validity of molecules produced by EQGAT-diff is slightly higher. Using SemlaFlow with 100 sampling steps corresponds to more than a 20x speedup over EQGAT-diff, and using 20 sampling steps corresponds to a two order-of-magnitude increase in sampling speed. Notably, however, molecules generated by EQGAT-diff have lower minimised energies than SemlaFlow, suggesting that their model is better at finding molecular conformations which have lower energy minima, while our model is better at producing lower strain energies.